# PAINT: Parallel-in-time Neural Twins for Dynamical System Reconstruction

### v1

@@ Abstract
Neural surrogates have shown great potential in simulating dynamical systems, while offering real-time capabilities. We envision Neural Twins as a progression of neural surrogates, aiming to create digital replicas of real systems. A neural twin consumes measurements at test time to update its state, thereby enabling context-specific decision-making. A critical property of neural twins is their ability to remain on-trajectory, i.e., to stay close to the true system state over time. We introduce Parallel-in-time Neural Twins (PAINT), an architecture-agnostic family of methods for modeling dynamical systems from measurements. PAINT trains a generative neural network to model the distribution of states parallel over time. At test time, states are predicted from measurements in a sliding window fashion. Our theoretical analysis shows that PAINT is on-trajectory, whereas autoregressive models generally are not. Empirically, we evaluate our method on a challenging two-dimensional turbulent fluid dynamics problem. The results demonstrate that PAINT stays on-trajectory and predicts system states from sparse measurements with high fidelity. These findings underscore PAINT’s potential for developing neural twins that stay on-trajectory, enabling more accurate state estimation and decision-making.  [^1] Code is available at https://github.com/ParticulateFlow/PAINT-Parallel_in_time_Neural_Twins
Refer to captionTrueAutoregressiveParallel-in-time0tht-httt+nt+n
Figure 1: Real world measurements allow for parallel-in-time models to stay on-trajectory, while autoregressive models may drift over time. Dashed lines represent sampled trajectories. Parallel-in-time models take a window of measurements and predict a distribution of subtrajectories which can be sampled and used for state and uncertainty estimation. Autoregressive models are prone to over-reliance on their autoregressive state (see Section 3.3) and typically depend on an initial state.
@@ 1 Introduction
Modern artificial intelligence is increasingly integrated into the modeling, monitoring, and control of complex dynamical systems. A key development in this domain are digital twins [23]: virtual models that run alongside a physical system, using live sensor data to update its predictions. Unlike standard simulations, a digital twin continuously adjusts its internal state to match the real system’s behavior, thereby enabling better monitoring and automated decision-making.
Neural surrogates emulate expensive, high-fidelity simulations and offer fast predictions while maintaining reasonable accuracy. Recent works have already demonstrated remarkable results on computational fluid dynamics (CFD) [44, 5, 70, 8], as well as on particle based systems as surrogate models for discrete element methods [7].
However, transforming neural surrogates into neural twins is a challenging endeavor. The current generation of neural surrogates can neither ingest nor adapt to real-time measurements during inference. Therefore, we ask the question: How to design neural surrogates that adapt at test time using real-time measurements?
We address this question by introducing Parallel-in-time Neural Twins (PAINT), a family of methods for reconstructing dynamical systems from measurements. At the core of PAINT is training a generative neural network to model the joint conditional probability of system states given the measurements. At inference, it reconstructs a sequence of states from measurements using a sliding-window approach. This makes our method interpretable and avoids any dependence on an initial condition.
Mathematically, we show that PAINT enables accurate state estimation from measurements due to being on-trajectory, i.e., staying close to the true trajectory even in presence of prediction errors. We also show that autoregressive models may possess this property under certain conditions (e.g. Kalman filters), but not in general. Our analysis also sheds light on a phenomenon we call over-reliance on the autoregressive state. It describes the problem of an autoregressive model over-relying on an off-trajectory state, whereas an ideal model would recognize such drift and reduce the state’s influence on the prediction.
@@
  • To the best of our knowledge, we are the first to train an on-trajectory neural twin for dynamical systems. Our method is interpretable, has an inherent measure of uncertainty and takes no prior assumptions on the initial state, therefore making it more widely applicable.
@@ 2 Background
Vectors and matrices are denoted in bold, e.g., $\bm{x}\in\mathbb{R}^{n}$ and $\bm{A}\in\mathbb{R}^{m\times n}$. General sequences such as $k,\dots,t]$ are denoted in brackets as $[k,t]$. We may also use brackets as subscripts over sequences, e.g., $\bm{x}_{[k,t]}=[\bm{x}_{k},\dots,\bm{x}_{t}]$. True or ground-truth probability distributions are denoted $p(\cdot)$, while functions or distributions parameterized by neural networks use subscripts $\theta$ for their parameters, e.g., $f_{\theta}$ or $p_{\theta}(\cdot)$.
2.1 Digital twins and data integration
Digital twins
as defined by Grieves [23] interact with an industrial process by (a) consuming real-time data and (b) taking decisions based on the current state of the system. Ozan et al. [56] marks a first end-to-end approach to incorporate and control a PDE with a learned policy. However as the state estimation is not interpretable, these models are hard to evaluate and can hardly be trusted for real-world production use cases. For a comprehensive introduction we refer to [65].
Bayesian filtering.
Updating model states by incorporating measurement data has a rich history, dating back to the seminal works on Wiener and Kalman filters [69, 34]. These types of Bayesian filters were extensively studied, leading to generalizations and extensions like the extended, unscented Kalman filters or particle methods. Bayesian filtering methods can be categorized by (1) Filtering: $p(\bm{x}_{t}\mid\bm{m}_{[0,t]})$, (2) Prediction: $p(\bm{x}_{t+n}\mid\bm{m}_{[0,t]})$, and (3) Smoothing: $p(\bm{x}_{t}\mid\bm{m}_{[0,T]})$, where $\bm{x}\in\mathbb{R}^{d_{x}},\bm{m}\in\mathbb{R}^{d_{m}},t\in[0,T]$ and $n\in\mathbb{N}$. We refer to Särkkä & Svensson [61] for a comprehensive overview on classical Bayesian filtering.
With the rise of deep learning in the mid-2010s, several works married Kalman or Bayesian filters with deep neural networks [40, 35, 37]. However, similar to the original Kalman filter these works usually rely on simplifying assumptions, on the model, noise or initial condition, which is in stark contrast to our method.
@@ 2.1 Digital twins and data integration
Prior art has explored flow reconstruction from sparse measurements. These measurements are usually assumed to come from Particle Tracking Velocimetry (PTV), probes in the walls or at informative locations. Some works used generative modeling techniques to reconstruct flow fields from sparse measurements for a single timestep without temporal context [24, 13, 53, 36, 25]. Similarly, Chakraborty et al. [11] trained an instantaneous super-resolution weather model. Several works used physics-informed neural networks (PINNs) [60] to reconstruct an instantaneous flow field from sparse measurements [12, 30, 66], however they lack the generative modeling which is necessary to rightfully model the distribution of possible solutions. Dang et al. [15] used spatiotemporal information for reconstruction, but they did not model the target trajectories probabilistically and therefore learn the theoretical mean of the conditional distribution. From a conceptual point of view, the closest to our concrete implementation of PAINT is Sun & Wang [64], however their construction based on Bayesian neural networks is unnecessarily complex, did only work on simple laminar flows and is not scalable to larger settings.
2.2 Scalable Neural Surrogates
Surrogates models.
A surrogate model, or simply a surrogate [20], is a simplified computational model designed to approximate a more intricate, computationally intensive system, such as a high-fidelity simulation or a physical experiment.
Transformers.
The application of transformers [67] as surrogate models has recently taken over, building on their established efficacy in diverse scientific domains, including protein folding [33, 1] and weather forecasting [9, 10]. These transformer-based neural surrogates are engineered to synthesize information across varying spatial locations and scales by leveraging attention mechanisms [67]. Transformer-based surrogates represent an extension of the neural operator framework [50, 51, 43, 42, 39], specifically by integrating self-attention, cross-attention, or perceiver blocks [32, 31]. Notable examples of these models include OFormer [45], Transolver [70], and (AB-)UPT [6, 8]. They have paved the way for scalable neural surrogate learning, scaling from a few thousand [44, 70, 6] to millions of mesh points [8].
Generative modeling of PDEs.
Over the past years, a powerful group of related generative modeling frameworks were introduced around the works of Diffusion [63, 27], Flow Matching [49, 46] and stochastic interpolants [2, 3, 22].
Generative models were used to model PDEs [62, 38]. Some works emphasize the beneficial properties of diffusion models for modeling high frequencies [48], leading to more accurate and physically plausible solutions. Our work differs, as prior works sample any trajectory $p(\bm{x}_{[0,t]})$, while we model $p(\bm{x}_{[0,t]}\mid\bm{m}_{[0,t]})$ and aim to stay on the real trajectory from measurements $\bm{m}_{[0,t]}$ (see Section 3.2).
@@ 2.3 Computational Fluid Dynamics
The dynamical system we will use in our experiments (see Section 4) is a 2D fluid dynamical system in the turbulent regime. Therefore, we give a brief introduction to computational fluid dynamics.
@@
where $\bm{u}$ denotes the velocity, $p$ denotes the pressure, $\mu$ denotes the viscosity and $\bm{f_{e}}$ denotes any kind of external body forces. These equations are complemented by appropriate boundary and initial conditions. This constitutes a system of equations of highly complex, second-order nonlinear PDEs with no known analytical solutions. Therefore, they are typically solved numerically.
@@
Numerical solutions of the Navier–Stokes equations involve a tradeoff between accuracy and computational cost. Direct numerical simulation (DNS) resolves the full flow field but is prohibitively expensive for most practical flows. Large-eddy simulation (LES) resolves only the largest, energy-containing scales, while Reynolds-averaged Navier–Stokes (RANS) approaches provide only mean flow fields under strong modeling assumptions [18, 21]. LES represents a compromise between fidelity and computational cost within this spectrum of methods and was therefore used for the experiments in this paper.
@@ 3.1 A neural twin framework for dynamical systems
DynamicalSystemSimulationNeural Twin(1) Inform(2) Training dataset(3) Measurements(4) Control
Figure 2: Interaction between a real-world dynamical system, simulations, and a neural twin. The workflow is structured in three phases: (1) The dynamical systems informs the design of the simulations. (2) Via the simulations, a diverse training dataset is generated, which is used for training the neural twin. (3) During inference, the neural twin integrates real-time measurement data to remain on-trajectory. (4) The state estimation of the neural twin can optionally be used for decision-making.
@@ 3.2 Parallel-in-time Neural Twins
As the number of timesteps may become very large, PAINT models the joint probability over a window (see Figure 1) of width $h$ and may predict $n$ timesteps into the future:
@@
PAINT is agnostic to the generative modeling framework and the neural network architecture. Common building blocks may be transformers, MLPs, CNNs [41].
@@ 3.3 Theoretical Considerations
where the autoregressive model learns the conditional probability $p_{\theta}(\bm{x}_{t}\mid\bm{x}_{t-1},\bm{m}_{[0,t]})$. Notice, this does not equal $p_{\theta}(\bm{x}_{t}\mid\bm{x}_{t-1},\bm{m}_{t})$, although e.g. Kalman filters use this form.
@@
Our analysis is leaned on the analysis of error growth of dynamical systems [55] and key observations made in [26]. We assume a fully differentiable neural network, which learns the probability distribution $p_{\theta}(\bm{x}_{t}||\bm{x}_{t-1},\bm{m}_{[0,t]})$. To sample from this probability distribution $\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t}||\bm{x}_{t-1},\bm{m}_{[0,t]})$ we rewrite the neural network to a deterministic function with a stochastic input $\eta_{t}$:
@@
The Jacobian is then defined as $\bm{J}_{t}\coloneqq\partial f_{\theta}/\partial\bm{x}_{t}$. In contrast to Orrell [55], this is a discrete random ordinary differential equation (RODE), i.e., an ODE with a random input at each timestep.
@@
The first-order Taylor approximation of a small input perturbation at $t=k$ is:
@@
This means, a deviation at timestep $k$ induces a deviation multiplied by the Jacobian product series at timestep $t$. This has also been experimentally explored by Hess et al. [26], who include visualizations of the Jacobian product series for chaotic dynamical systems.
@@
In the following we will formalize the notions of a model being “on-trajectory”. For this, we rely on Assumption 1, which states that the predictive relevance of past measurements decays within a finite window $h$.
Assumption 1.
Temporal decay of measurement information: For the given dynamical system, there is a finite window size $h\in\mathbb{N}$, such that almost all information from past measurements comes from measurements within this window. Formally, for all $\epsilon>0$ and all $t\in\mathbb{N}$ there exists an $h\ll t$ such that $||p(\bm{x}_{t}\mid\bm{m}_{[t-h,t]})-p(\bm{x}_{t}\mid\bm{m}_{[0,t]})||<\epsilon$.
@@
Definition 1.
On-trajectory: Assuming unbounded model size, compute and data, a model is on-trajectory if the modeled distribution is arbitrarily close to the true distribution. Formally, for all $t\in\mathbb{N}$ and any $\epsilon>0$ it holds that: $||p_{\theta}(\bm{x}_{t}\mid\bm{m}_{[0,t]})-p(\bm{x}_{t}\mid\bm{m}_{[0,t]})||<\epsilon$.
Parallel-in-time models are on-trajectory under Assumption 1 .
We derive this in Appendix B.1. Intuitively, for each $\epsilon$ we can choose the window $h$ large enough to guarantee arbitrary closeness. In contrast, autoregressive models are in general not on-trajectory. As a counterexample, consider an underlying chaotic system where the model ignores the measurements (for further discussion, see Hess et al. [26] and Appendix B.2).
Over-reliance on the autoregressive state.
The above analysis highlights a fundamental problem we call over-reliance on the autoregressive state. Intuitively, it states that an autoregressive state plays an ambiguous role. If the state is on-trajectory, it helps the model as it provides useful information for the next state. If the state is off-trajectory, the model may still make use of it for prediction and drift off further. In principle, the only way for the model to correct an off-trajectory autoregressive state is with informative measurements. This excludes edge cases, e.g., where the model converges to a global steady state.
Ideally, a model would detect when the autoregressive state drifts off and decrease its influence on the prediction. Depending on the informativeness of the measurements, it could increase its reliance on measurement information or capture the inherent uncertainty in a different way.
Over-reliance on autoregressive state is also influenced by the training strategy. Autoregressive models are usually trained with teacher-forcing, where the ground-truth previous state $\bm{x}_{t}$ is used to predict the next state $\bm{x}_{t+1}$. As these states are highly correlated, the model might learn to over-rely on $\bm{x}_{t}$ and under-rely on $\bm{m}_{t+1}$. This could potentially be mitigated by training techniques such as generalized teacher forcing [26] or unrolled training [38].

### v2

@@ Abstract
Neural surrogates have shown great potential in simulating dynamical systems, while offering real-time capabilities. We envision Neural Twins as a progression of neural surrogates, aiming to create digital replicas of real systems. A neural twin consumes measurements at test time to update its state, thereby enabling context-specific decision-making. We argue, that a critical property of neural twins is their ability to remain on-trajectory, i.e., to stay close to the true system state over time. We introduce Parallel-in-time Neural Twins (PAINT), an architecture-agnostic family of methods for modeling dynamical systems from measurements. PAINT trains a generative neural network to model the distribution of states in parallel over time. At test time, states are predicted from measurements in a sliding window fashion. Our theoretical analysis shows that PAINT is on-trajectory, whereas autoregressive models generally are not. Empirically, we evaluate our method on a challenging two-dimensional turbulent fluid dynamics problem. The results demonstrate that PAINT stays on-trajectory and predicts system states from sparse measurements with high fidelity. These findings underscore PAINT’s potential for developing neural twins that stay on-trajectory, enabling more accurate state estimation and decision-making.  [^1] Code is available at https://github.com/ParticulateFlow/PAINT-Parallel_in_time_Neural_Twins
Refer to caption TruePAINTAutoregressive0tht-httt+nt+n
Figure 1: Conceptual depiction of trajectories from the ground truth (blue), our parallel-in-time method PAINT (green) and a naive autoregressive model (orange). Dashed lines represent sampled trajectories from the probability distribution (solid lines) given by the respective model.
The models ingest real world measurements (not depicted) at test time to reconstruct the true trajectory. Autoregressive models predict the probability of the next state given an autoregressive state additional to real world measurements. They typically depend on an initial state and are prone to drift off due to over-reliance on their autoregressive state (see Section 3.4). In contrast, parallel-in-time models are “on-trajectory” and do not depend on an initial state. They take a window of measurements and predict a distribution of subtrajectories which can be sampled and used for state and uncertainty estimation.
@@ 1 Introduction
Modern artificial intelligence is increasingly integrated into the modeling, monitoring, and control of complex dynamical systems. A key development in this domain are digital twins (Grieves, 2014): virtual models that run alongside a physical system, using live sensor data to update its predictions. Unlike standard simulations, a digital twin continuously adjusts its internal state to match the real system’s behavior, thereby enabling better monitoring and automated decision-making.
Neural surrogates are neural network based predictive models which emulate expensive, high-fidelity simulations and offer fast predictions while maintaining reasonable accuracy. Recent works have already demonstrated remarkable results on computational fluid dynamics (CFD) (Li et al., 2022; Alkin et al., 2024a; Wu et al., 2024; Alkin et al., 2025), as well as on particle based systems as surrogate models for discrete element methods (Alkin et al., 2024c).
However, transforming neural surrogates into neural twins is a challenging endeavor. The current generation of neural surrogates can neither ingest nor adapt to real-time measurements during inference. Therefore, we asked the question: How to design neural surrogates that adapt at test time using real-time measurements?
We do not claim to solve this problem entirely. However, we present insights and a new approach that advances the discussion. We introduce Parallel-in-time Neural Twins (PAINT), a family of methods for reconstructing dynamical systems from measurements. At the core of PAINT is training a generative neural network to model the joint conditional probability of system states given the measurements. At inference, it reconstructs a sequence of states from measurements using a sliding-window approach. This makes our method interpretable and avoids any dependence on an initial condition, which constitutes an easily overlooked component in autoregressive models.
Mathematically, we show that PAINT enables accurate state estimation from measurements due to being on-trajectory, i.e., staying close to the true trajectory even in presence of prediction errors. We also show that autoregressive models may possess this property under certain conditions, but not in general. Our analysis also sheds light on a phenomenon we call over-reliance on the autoregressive state. It describes the problem of an autoregressive model over-relying on an off-trajectory state, whereas an ideal model would recognize such drift and reduce the state’s influence on the prediction.
@@
  • To the best of our knowledge, we are the first to train an on-trajectory neural twin on a high-dimensional fluid-dynamics problem. Our method is interpretable, has an inherent measure of uncertainty and makes no prior assumptions about the initial state.
@@ 2 Background and related work
Vectors and matrices are denoted in bold, e.g., $\bm{x}\in\mathbb{R}^{n}$ and $\bm{A}\in\mathbb{R}^{m\times n}$. General sequences such as $k,\dots,t]$ are denoted in brackets as $[k,t]$. We may also use brackets as subscripts over sequences, e.g., $\bm{x}_{[k,t]}=[\bm{x}_{k},\dots,\bm{x}_{t}]$. True or ground-truth probability distributions are denoted $p(\cdot)$, while the probability of a random variable $\bm{X}$ taking a specific value $\bm{x}$ is written as $p(\bm{X}=\bm{x})$ or, in shorthand, as $p(\bm{x})$. Functions or distributions parameterized by neural networks use subscripts $\theta$ for their parameters, e.g., $f_{\theta}$ or $p_{\theta}(\cdot)$.
2.1 Scalable probabilitic neural surrogates
Scalable neural surrogates.
A surrogate model, or simply a surrogate (Forrester et al., 2008), is a simplified computational model designed to approximate a more intricate, computationally intensive system, such as a high-fidelity simulation or a physical experiment. Modern neural surrogates based on deep neural networks rely heavily on the transformer architecture (Vaswani et al., 2017). They have been successfully scaled from a few thousand (Li et al., 2022; Wu et al., 2024; Alkin et al., 2024b) to millions of input dimensions (Alkin et al., 2025). Often these architectures are combined with convolutional layers (LeCun et al., 1989; Kipf, 2016) as learned encoders and decoders (Kohl et al., 2023; Alkin et al., 2025).
Generative modeling of PDEs.
Over the past years, a powerful group of related generative modeling frameworks were introduced around the works of Diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020), Flow Matching (Liu et al., 2022; Lipman et al., 2022) and stochastic interpolants (Albergo & Vanden-Eijnden, 2022; Albergo et al., 2023a; Gao et al., 2024). These frameworks have set new state-of-the-art results when modeling PDEs (Sestak et al., 2025; Kohl et al., 2023). Additionally diffusion models seem to possess beneficial properties for modeling high frequencies (Lippe et al., 2023), leading to more accurate and physically plausible solutions.
2.2 Reconstruction and control of dynamical systems
Bayesian inference.
Updating model states by incorporating measurement data is known as bayesian inference (Särkkä & Svensson, 2023). Bayesian inference methods can be broadly categorized into (1) Filtering: $p(\bm{x}_{t}\mid\bm{m}_{[0,t]})$, (2) Prediction: $p(\bm{x}_{t+n}\mid\bm{m}_{[0,t]})$, and (3) Smoothing: $p(\bm{x}_{t}\mid\bm{m}_{[0,T]})$, where $\bm{x}\in\mathbb{R}^{d_{x}},\bm{m}\in\mathbb{R}^{d_{m}},t\in[0,T]$ and $n\in\mathbb{N}$. Classical approaches to bayesian inference take simplifying assumptions and usually arrive at analytical solutions. These approaches date back to the seminal works on Wiener and Kalman filters (Wiener, 1949; Kalman, 1960). However, neither classical approaches nor hybrid approaches using neural networks (Krishnan et al., 2015; Karl et al., 2016; Kim et al., 2025) have been successfully trained on tens of thousands or more input dimensions. They are therefore impractical for our use case.
@@ 2.2 Reconstruction and control of dynamical systems
Prior art has explored flow reconstruction from sparse measurements. These measurements are usually assumed to come from Particle Tracking Velocimetry (PTV), probes in the walls or at informative locations.
PAINT differentiates and/or generalizes most prior works by (1) learning a distribution of states instead of point estimates (Dang et al., 2024; Chaurasia & Chakraborty, 2024; Hosseini & Shiri, 2024; Toscano et al., 2024) or (2) being scalable and end-to-end trainable (Sun & Wang, 2020). or (3) using temporal additionally to spatial information (Güemes et al., 2022; Cuéllar et al., 2024; Oommen et al., 2025; Kim et al., 2021; Hemant Parikh et al., 2025).
Some works (Li et al., 2024; Du et al., 2024; Amorós-Trepat et al., 2026) used diffusion posterior sampling (Jalal et al., 2021; Chung et al., 2022) to reconstruct flow fields from measurements based on a prior diffusion model. However, the quality of the reconstruction still falls short of conditional diffusion models (Gupta et al., 2024).
Control of dynamical systems.
Instead of reconstructing the dynamical system from measurements, prior work controlled the dynamical system directly from the measurements using reinforcement learning (Ozan et al., 2025; Lagemann et al., 2025). While these are promising directions, the lack of interpretability is a major obstacle for a broader adoption and a limitation regarding applicability.
@@ 2.3 Computational Fluid Dynamics
The dynamical system we will use in our experiments (see Section 4.1) is a 2D fluid dynamical system in the turbulent regime. Therefore, we give a brief introduction to computational fluid dynamics.
@@
where $\bm{u}$ denotes the velocity, $p$ denotes the pressure, $\mu$ denotes the viscosity and $\bm{f_{e}}$ denotes any kind of external body forces. These equations are complemented by appropriate boundary and initial conditions. This constitutes a system of equations of highly complex, second-order nonlinear PDEs with no known general analytical solution. Therefore, they are typically solved numerically.
@@
Numerical solutions of the Navier–Stokes equations involve a tradeoff between accuracy and computational cost. Direct numerical simulation (DNS) resolves the full flow field but is prohibitively expensive for most practical flows. Large-eddy simulation (LES) resolves only the largest, energy-containing scales, while Reynolds-averaged Navier–Stokes (RANS) approaches provide only mean flow fields under strong modeling assumptions (Davidson, 2015; Fröhlich, 2006). LES represents a compromise between fidelity and computational cost within this spectrum of methods and was therefore used for the experiments in this paper.
@@ 3 Method
DynamicalSystemSimulationdatasetNeural Twin(1) Inform(2) Train(3) Measurements(4) Control
Figure 2: Interaction between a real-world dynamical system, the simulations dataset and a neural twin. The training phase is depicted in blue, the inference (test) phase in red and nodes involved in both phases in black. Dashed lines indicate optional steps. The workflow is structured in three phases: (1) The dynamical systems informs the design of the simulations and a diverse training dataset is generated. (2) The generated dataset is used for training the neural twin. (3) During inference, the neural twin integrates real-time measurement data to remain on-trajectory. (4) The state estimation of the neural twin can optionally be used for decision-making.
@@ 3.2 Parallel-in-time Neural Twins
As the number of timesteps may become very large, PAINT models the joint probability over a window (see Figure 1) of width $h$ and may optionally predict $n$ timesteps into the future:
@@
PAINT is agnostic to the generative modeling framework and the neural network architecture. Common building blocks may be transformers, MLPs or CNNs (LeCun et al., 1989).
@@ 3.3 Theoretical Considerations
$p(\bm{x}_{[0,t]}\mid\bm{m}_{[0,t]})=p(\bm{x}_{0}\mid\bm{m}_{[0,t]})\cdot\prod_{i=1}^{t}p(\bm{x}_{i}\mid\bm{x}_{i-1},\bm{m}_{[0,t]}),$ (3)
where the autoregressive model learns the conditional probability $p_{\theta}(\bm{x}_{t}\mid\bm{x}_{t-1},\bm{m}_{[0,t]})$. However, in practice, due to finite compute but potentially arbitrarily long sequences, a shorter measurement window of size $h\in\mathbb{N}$ is used as follows
$p_{\theta}(\bm{x}_{t}\mid\bm{x}_{[t-k,t-1]},\bm{m}_{[t-h,t]})$ (4)
@@
Our analysis is leaned on the analysis of error growth of dynamical systems (Orrell, 2005) and key observations made in (Hess et al., 2023). We assume a fully differentiable neural network, which learns the probability distribution $p_{\theta}(\bm{x}_{t}||\bm{x}_{t-1},\bm{m}_{[0,t]})$. To sample from this probability distribution $\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t}||\bm{x}_{t-1},\bm{m}_{[0,t]})$ we rewrite the neural network to a deterministic function with a stochastic input $\eta_{t}$:
@@
The Jacobian is then defined as $\bm{J}_{t}\coloneqq\partial f_{\theta}/\partial\bm{x}_{t}$. In contrast to Orrell (2005), this is a discrete random ordinary differential equation (RODE), i.e., an ODE with a random input at each timestep.
@@
The first-order Taylor approximation of a small input perturbation at timestep $k<t$ is:
@@
This means, a deviation at timestep $k$ induces a deviation multiplied by the Jacobian product series at timestep $t$. This has also been experimentally explored by Hess et al. (2023), who include visualizations of the Jacobian product series for chaotic dynamical systems.
@@
In the following we will formalize the notion of a model being “on-trajectory”. For this, we rely on Assumption 3.1, which states that the predictive relevance of past measurements decays within a finite window $h$.
Assumption 3.1.
Temporal decay of measurement information: For the given dynamical system, there is a finite window size $h\in\mathbb{N}$, such that almost all information from past measurements comes from measurements within this window. Formally, for all $\epsilon>0$ and all $t\in\mathbb{N}$ there exists a $h\ll t$ such that $||p(\bm{x}_{t}\mid\bm{m}_{[t-h,t]})-p(\bm{x}_{t}\mid\bm{m}_{[0,t]})||<\epsilon$.
@@
Definition 3.2.
On-trajectory: Assuming unbounded model size, compute and data, a model is on-trajectory if the modeled distribution is arbitrarily close to the true distribution. Formally, for all $\epsilon>0$ and all $t\in\mathbb{N}$, $\bm{x}\in\bm{\mathcal{X}}$ there is model, data and compute such that: $||p_{\theta}(\bm{X}_{t}=\bm{x}\mid\bm{m}_{[0,t]})-p(\bm{X}_{t}=\bm{x}\mid\bm{m}_{[0,t]})||<\epsilon$.
Parallel-in-time models are on-trajectory under Assumption 3.1 .
We derive this in Appendix A.1. Intuitively, for each $\epsilon$ we can choose the window $h$ large enough to guarantee arbitrary closeness. In contrast, autoregressive models are in general not on-trajectory. As a constructed counterexample, consider an underlying chaotic system where the model learns to ignore the measurements (for further discussion, see Hess et al. (2023) and Appendix A.2).
3.4 Over-reliance on the autoregressive state
The analysis in Section 3.3 highlights a fundamental problem we call over-reliance on the autoregressive state. Intuitively, it states that for a naive autoregressive model the autoregressive state plays an ambiguous role: If the state is on-trajectory, it helps the model as it provides useful information for the next state. If the state is off-trajectory, the model may still make use of it for prediction and drift off further.
Measurement informativeness and uncertainty modeling.
In principle, the only way for the model to correct an off-trajectory autoregressive state is with informative measurements (this excludes edge cases, e.g., where the model converges to a global steady state). Ideally, a model would detect when the autoregressive state drifts off and decrease its influence on the prediction. This can theoretically be achieved by a model which models the state uncertainty explicitly. However reliable uncertainty modeling for high-dimensional state spaces and long trajectories is a challenging open research problem.
Training paradigm.
Over-reliance on autoregressive state is also influenced by the training strategy. Autoregressive models are usually trained with teacher-forcing, where the ground-truth previous state $\bm{x}_{t}$ is used to predict the next state $\bm{x}_{t+1}$. As these states are highly correlated, the model might learn to over-rely on $\bm{x}_{t}$ and under-rely on $\bm{m}_{t+1}$. This could potentially be mitigated by training techniques such as generalized teacher forcing (Hess et al., 2023) or unrolled training (Kohl et al., 2023).
Connection to Jacobian.